What are the coordinates of the turning point of the graph with the equation y=3x^2+6x-1?
Answer by Alan3354(69443)

What are the coordinates of the turning point of the graph with the equation y=3x^2+6x-1?
-------------
The vertex is at x = -b/2a = -6/6 = -1
y = -4
-->(-1,-4) is the vertex, a minimum.
Is that what you mean by "turning point?" There is no point of inflection, the 2nd derivative is 6, no sign change.